an electron is sitting 10 cm away from a small object of charge -5nc it is let go and move away from the charge object when it reaches a distance of 15 cm away. find its speed
1
Expert's answer
2019-10-29T11:03:29-0400

Using the energy conservation principle:


"\frac{kq_1q_2}{r_0}=\frac{kq_1q_2}{r_1}+\frac{mv^2}{2}"


So:


"v=\sqrt{\frac{2}{m}(\frac{kq_1q_2}{r_0}-\frac{kq_1q_2}{r_1})}"

"v=\sqrt{\frac{2}{9.1*10^{-31}}(9*10^9*1.6*10^{-19}*5*10^{-9}(\frac{1}{0.1}-\frac{1}{0.15}))}"

"v=7.26*10^6 m\/s"